# Institutionalizing Folk Theories of Algorithms: How Multi-Channel Networks (MCNs) Govern Algorithmic Labor in Chinese Live-Streaming Industry

### v1

@@ Abstract.
As algorithmic systems increasingly structure platform labor, workers often rely on informal “folk theories”, experience-based beliefs about how algorithms work, to navigate opaque and unstable algorithmic environments. Prior research has largely treated these theories as bottom-up, peer-driven strategies for coping with algorithmic opacity and uncertainty. In this study, we shift analytical attention to intermediary organizations and examine how folk theories of algorithms can be institutionally constructed and operationalized by those organizations as tools of labor management. Drawing on nine months of ethnographic fieldwork and 37 interviews with live-streamers and staff at Multi-Channel Networks (MCNs) in China, we show that MCNs develop and circulate dual algorithmic theories: internally, they acknowledge the volatility of platform systems and adopt probabilistic strategies to manage risk; externally, they promote simplified, prescriptive theories portraying the algorithm as transparent, fair, and responsive to individual effort. They have further operationalize those folk theories for labor management, encouraging streamers to self-discipline and invest in equipment, training, and routines, while absolving MCNs of accountability. We contribute to CSCW and platform labor literature by demonstrating how informal algorithmic knowledge, once institutionalized, can become infrastructures of soft control – shaping not only how workers interpret platform algorithms, but also how their labor is structured, moralized and governed.
Algorithm; Platform; Folk Theories of Algorithms; Live-streaming; MCNs; China; Ethnography
@@
This work was completed during Xianzhe Fan’s visiting research at Carnegie Mellon University.
@@ 1. Introduction
As algorithmic systems become increasingly central to digital labor, workers across domains, ride-hailing, content moderation, e-commerce, and live-streaming, are compelled to navigate opaque, volatile, and often inscrutable mechanisms of evaluation and visibility. Algorithms do not simply mediate tasks; they embed managerial logics into platform infrastructure, determining what is seen, who is rewarded, and how performance is measured (Wood et al., 2019; Bucher, 2018; Kellogg et al., 2020). For workers, this opacity creates a pervasive sense of precarity: their success of the work appears contingent not only on effort or skill, but on deciphering hidden systems that govern algorithmic recognition. In response, scholars in HCI, CSCW, and media studies have increasingly turned their attention to the interpretive practices workers develop to make sense of algorithmic environments. A key concept emerging from this line of research is folk theories of algorithms—informal beliefs based on personal experience about how algorithms function (DeVito et al., 2017; Xiao et al., 2024b; Ytre-Arne and Moe, 2021). These theories allow workers to construct actionable understandings of platform behavior, guiding decisions about when, how, and what to produce. Prior work has richly documented how workers share folk theories in peer networks (McDaid et al., 2023; Yao et al., 2021), use them to rationalize inconsistent outcomes (Christin, 2020; Poon et al., 2021), or deploy them to strategically align with perceived algorithmic preferences (Bonini and Treré, 2024; Lei, 2021). More recent scholarship also examines how these theories are not developed in isolation but are shaped collectively within communities (Xiao et al., 2024b; Zhang et al., 2023). However, this growing body of work continues to frame folk theories primarily as bottom-up epistemic practices—emergent, community-driven, and reactive. What remains underexplored is how such theories can be deliberately organized, curated, and operationalized by third-party institutions. That is, how folk theories shift from informal peer-to-peer artifacts to formalized narratives embedded in managerial routines, spatial arrangements, and incentive structures.
@@
This paper addresses this gap by examining the role of Multi-Channel Networks (MCNs) in China’s live-streaming industry. MCNs are intermediary organizations that operate between individual streamers and platform companies. They help streamers get started, improve their content, build their personal brand, and make money, but also often control what streamers produce and how they work (Zhang et al., 2024; Zhang and Tong, 2024; Liang and Ji, 2024). While MCNs have been widely studied in western contexts such as YouTube as service providers that support monetization, branding, and creator management (Siciliano, 2023; Gardner and Lehnert, 2016), these research often overlook their epistemic role, especially in algorithm interpretation. We advance this literature by theorizing MCNs not just as economic intermediaries, but as organizational actors that actively interpret and institutionalize algorithmic knowledge. This shift foregrounds how algorithmic sensemaking is produced and circulated at scale. China’s live-streaming industry then offers a particularly generative setting for this analysis. Unlike more entertainment-driven and part-time creator cultures elsewhere, Chinese live-streaming is highly industrialized and labor-intensive (Ye et al., 2023; Sun et al., 2023), with MCNs exerting deep control over labor and visibility. Despite their waning influence on Western markets (Alexander, 2018), MCNs remain dominant in China, shaping how streamers understand and respond to algorithmic systems. Given the generally low algorithmic literacy of streamers (Xiao et al., 2024a; Tang et al., 2022), MCNs possess disproportionate epistemic authority, determining not just what is knowable about algorithms, but who gets to know it and how.
@@
In this study, we focus on how Chinese MCNs develop and deploy folk theories of algorithms organizationally for labor management in live-streaming industry. Drawing on nine months of ethnographic fieldwork and 37 interviews with MCN staff and streamers, we ask:
  • RQ1: How do MCNs interpret and shape streamers’ understanding of algorithms on Chinese live-streaming platforms?
@@
  • RQ2: How are these interpretations translated into everyday organizational practices for labor management?
Rather than treating folk theories of algorithms as bottom-up sensemaking tools, our findings reveal how they are deliberately constructed, circulated, and embedded by intermediary organizations (MCNs) to govern labor under algorithmic opacity. Specifically, we find that MCNs construct dual folk theories of algorithms: internally, they acknowledge the instability and opacity of platform algorithms, using this uncertainty to manage organizational risk; externally, they promote motivational narratives that frame the algorithm as fair, responsive, and effort-driven, shifting the burden of performance onto individual streamers. These folk theories are institutionalized through moral discourse, performance rankings, aesthetic standards, spatial control, and financial dependencies, making speculative knowledge of algorithms feel actionable and enforceable.
This study offers three key contributions to the CSCW/HCI literature. First, it re-theorizes folk algorithmic knowledge as institutionalized epistemic infrastructure, not merely emergent and experiential, but strategically produced and maintained within organizations. Second, it conceptualizes these folk theories as tools of algorithmic labor management, enabling MCNs to govern workers through belief, aspiration, and moral obligation. Third, it provides design and policy implications for CSCW and HCI, urging system designers and policy makers to account for epistemic authority that mediates algorithmic work.
@@ 2. Background and Related Work
This section reviews three strands of literature that form the foundation of our study. First, we examine research on folk theories of algorithms (subsection 2.1). Second, we analyze the shift from direct algorithmic management (e.g., task allocation or performance metrics) to softer and more affective forms of control (subsection 2.2). Third, we turn to literature on Multi-Channel Networks (MCNs) and surrounding studies on the live-streaming industry (subsection 2.3).
2.1. Folk Theories of Algorithms
In this section, we examine two interrelated dimensions around algorithms. First, we explore how workers, both individually and collectively, construct informal algorithmic knowledge to adapt and survive amid opacity and volatility (subsubsection 2.1.1). Second, we investigate how organizations actively shape and institutionalize folk theories to align worker behavior with algorithmic logics and organizational goals (subsubsection 2.1.2).
2.1.1. From Individual to Collective Folk Theories of Algorithms
As algorithmic systems have grown increasingly opaque, unstable, and consequential, users and workers have developed informal interpretive frameworks, often termed folk theories of algorithms, to perceive, understand and make sense of how these systems function (DeVito et al., 2017; Bucher, 2018; Dogruel, 2021; Huang et al., 2022). Rooted in everyday experiences rather than technical knowledge, such folk theories reflect individuals’ efforts to explain, predict, and act upon algorithmic outcomes in the absence of transparent information, often drawing on trial and error, peer advice, or rumor  (Ytre-Arne and Moe, 2021; Siles et al., 2020), while most studies frame folk theories as enabling users to preserve a sense of comprehension and control despite the opacity of algorithmic systems  (Eslami et al., 2016; Rader and Gray, 2015).
In the context of platform labor, folk theories also serve as practical guides among workers to navigating algorithmic systems, such as those that determine content visibility or task allocation (McDaid et al., 2023; Bishop, 2020, 2019; DeVito, 2022). Workers on platforms like YouTube  (Bishop, 2020, 2019), TikTok  (DeVito, 2022), or Uber  (Cameron et al., 2021b, a; Cameron, 2020) often develop personalized work patterns from folk theories that shape their daily routines, even if these are technically inaccurate. Some scholars observed that ride-share drivers, confronted with opaque algorithmic decision-making, developed personalized explanations and workaround strategies to navigate the system, which were frequently discussed and refined within online peer communities  (Tarafdar et al., 2023; Ramesh et al., 2023; Lee et al., 2015; Rosenblat, 2018). In such settings, folk theories are rarely the product of isolated reasoning. Instead, they often emerge through collective construction, a process whereby individual experiences, suspicions, and partial observations are aggregated, debated, and iteratively reinforced within worker networks  (Xiao et al., 2024b; DeVito et al., 2017; Bonini and Treré, 2024; Karizat et al., 2021). Over time, these collaboratively formed beliefs can crystallize into shared logics that inform daily labor practices, regardless of their technical accuracy. These theories are not static; they circulate dynamically across online forums, social media groups, and informal conversations, adapting to shifts in platform policies or worker experiences  (DeVito, 2021; Bonini and Treré, 2024).
Thus, folk theories of algorithms operate as collective coping mechanisms, socially constructed and socially sustained forms of knowledge that make opaque algorithmic systems actionable and survivable in practice  (Xiao et al., 2024b). Yet, their influence extends beyond individual adaptation: as they circulate, stabilize, and sediment into routine, these theories begin to shape the very structure of platform work itself. They define what counts as “smart” strategies, legitimate routines, or risky behaviors under algorithmic governance, thereby producing normative expectations that guide how workers plan, perform, and assess their labor. Over time, these collectively maintained theories can crystallize into informal occupational scripts—shared understandings about how to succeed, what to avoid, and how to interpret algorithmic feedback  (Cameron, 2024; Meijerink and Bondarouk, 2023). Such scripts inform not only individual choices but also broader cultural logics within platform economies, reinforcing ideals of hustle, responsiveness, or creative optimization  (Xiao et al., 2024b). In this sense, folk theories are not merely responses to uncertainty; they are productive forces that help structure platform labor itself, embedding behavioral norms, emotional orientations, and survival tactics into the everyday life of algorithmic work  (Kellogg et al., 2020). To fully grasp the governance potential of folk theories, we must turn to the organizational and platform forces that selectively shape, formalize, and instrumentalize these beliefs.
2.1.2. From Interpretive Folk Theories to Organizational Control
While much of the existing literature conceptualizes folk theories as grassroots knowledge emerging from lived experience, a growing body of research highlights that such theories can also be constructed, curated, and institutionalized by organizations (Siciliano, 2023; Zhang and Tong, 2024). An emerging research direction asks how platform organizations, managers, or intermediary institutions actively shape, amplify, or exploit these user beliefs as part of their strategic governance toolkit  (Cornils, 2020; Randhawa et al., 2017; Ticona, 2022; Xiao et al., 2024b).
Platforms play a foundational role in shaping user beliefs through deliberate ambiguity, selective disclosure, and interface nudges. For example, Uber, the platform’s vague deactivation policies and gamified app interface led many drivers to internalize the belief that low acceptance rates would lead to penalties, even though this was never officially confirmed (Rosenblat and Stark, 2016; Heiland, 2025). Uber did little to dispel this belief, and it arguably served the company’s interest by promoting behavioral compliance. Similarly, Ma and Kou’s study on YouTube demonstrates how opaque demonetization practices forced creators to develop their own moderation heuristics, such as avoiding controversial keywords or altering content format, to appease the perceived preferences of the monetization algorithm (Ma and Kou, 2021). In both cases, the platforms shaped the contours of folk theories by withholding definitive explanations while maintaining algorithmic authority, thereby offloading interpretive labor onto users.
As Rosenblat and Stark argue in their study of Uber, the asymmetries of information and power embedded in the platform’s technical and communicative infrastructure are central to its ability to manage workers without formal employment structures. In their view, algorithmic systems are not only functional tools but also rhetorical devices that legitimize corporate control while disempowering labor  (Rosenblat and Stark, 2016). This discursive and infrastructural opacity enables platforms to offload interpretive labor onto workers, who in turn develop folk theories that align their behavior with the platform’s operational goals, even when these theories are based on speculation. These insights form the theoretical foundation of our study: that sometimes users’ understandings of algorithms can be deliberately shaped and directed by institutional actors.
What remains underexamined, however, is how these processes unfold not within platforms themselves, but through third-party actors that mediate between platforms and workers. If platforms benefit from maintaining algorithmic ambiguity, what role do intermediary organizations play in translating, amplifying, or formalizing that ambiguity? Recent studies have begun to explore this intermediary layer. For instance, Siciliano demonstrates how YouTube-based MCNs provide training and support to creators based on speculative algorithmic interpretations, reinforcing platform logics under the guise of optimization  (Siciliano, 2023). Xiao et al. examine how fan clubs as a force beyond platforms constructed an elaborate rhetorical discourse around algorithmic operations, along with concrete manipulation tutorials, to mobilize millions of ordinary users into collective action aimed at influencing recommendation algorithms  (Xiao et al., 2024b). These works on how third-party actors beyond platforms act in algorithmic labor governance and bring attention to the sociotechnical infrastructures through which folk theories circulates. However, we still know little about how these institutions embed folk theories into material and contractual infrastructures that shape labor discipline. To address these questions, our study investigates how algorithmic uncertainty is interpreted, standardized, and embedded into institutional practices, not merely as speculation, but as an instrument of labor governance in Chinese live-streaming industry.

### v2

# Constructing Algorithmic Authority: How Multi-Channel Networks (MCNs) Govern Live-Streaming Labor in China

@@ Abstract.
This study examines the discursive construction of algorithms and its role in labor management in Chinese live-streaming industry by focusing on how intermediary organizations (Multi-Channel Networks, MCNs) actively construct, stabilize, and deploy particular interpretations of platform algorithms as instruments of labor management. Drawing on a nine-month ethnographic fieldwork and 44 interviews with live-streamers, former live-streamers, and MCN staff, we examine how MCNs produce and circulate structured interpretations of platform algorithms across organizational settings. We show that MCNs articulate two asymmetric yet interconnected forms of algorithmic interpretations. Internally, MCNs managers approach algorithms as volatile and uncertain systems and adopt probabilistic strategies to manage performance and risk. Externally, in interactions with streamers, MCNs circulate simplified and prescriptive algorithmic narratives that frame platform systems as transparent, fair, and responsive to individual effort. These organizationally produced algorithmic interpretations are embedded into training materials, live-streaming performance metrics, and everyday management practices. Through these mechanisms, streamers internalize responsibility for outcomes, intensify self-discipline, and increase investments in equipment, performing skills, and routines to maintain streamer-audience relationship, while accountability for unpredictable outcomes is increasingly shifted away from managers and platforms. This study contributes to CSCW and platform labor research by demonstrating how discursively constructed algorithmic knowledge can function as an intermediary infrastructure of soft control, shaping how platform labor is regulated, moralized, and governed in practice.
Algorithm; Platform; Discursive Construction of Algorithms; Live-streaming; Multi-Channel Networks (MCNs); China; Ethnography
@@ 1. Introduction
As algorithmic systems have increasingly become central to digital labor, workers across domains such as ride-hailing, e-commerce, and live-streaming are required to operate within opaque, volatile, and hard-to-interpret systems of performance and evaluation. Algorithms embed managerial logic into platform infrastructures by shaping what becomes visible, how performance is assessed, and which forms of labor receive recognition or reward (Wood et al., 2019; Bucher, 2018; Kellogg et al., 2020). For workers, everyday labor unfolds under conditions of uncertainty, where outcomes appear closely tied to shifting algorithmic signals. These conditions give rise to sustained interpretive work as workers seek to make sense of how algorithmic systems operate and how their activity may be evaluated.
Scholars in HCI, CSCW, and digital studies have documented how workers develop practical understandings of algorithmic systems through experience. Across this literature, these interpretive processes have been theorized using concepts such as algorithmic imaginaries (Bucher, 2018), mental models (Ngo et al., 2020), and folk theories of algorithms (DeVito et al., 2017). Empirically, this body of research shows how workers exchange interpretations within peer networks (McDaid et al., 2023; Yao et al., 2021), use shared experiences and explanations to rationalize algorithmic outcomes (Christin, 2020; Poon et al., 2021), and align their practices with perceived algorithmic preferences (Bonini and Treré, 2024; Lei, 2021). Recent work further highlights the collective nature of these interpretive processes, demonstrating how understandings of algorithms circulate within communities and become stabilized through repeated interaction (Xiao et al., 2024b; Zhang et al., 2023). Together, these studies establish algorithmic interpretation as a central component of contemporary platform labor.
Moreover, prior work also treats algorithms as discursively constructed objects, examining how institutions articulate and frame particular understandings of algorithmic systems through their explanations (Ananny and Crawford, 2018; Cellard, 2022; Schulz, 2023). Lin ((2025)) introduces the term “discursive construction of algorithms” to capture the discursively produced character of algorithmic interpretations. Lin ((2025)) shows that Douyin’s official accounts foreground algorithmic fairness and effort-responsiveness while downplaying uncertainty, thereby reducing users’ suspicion of unfairness and aligning user behavior with the platform’s profit model. While this perspective clarifies how platforms shape users’ orientations toward algorithmic systems, we still know much less about how such discursive constructions become integrated into labor management through intermediary organizations.
Within platformized labor markets, intermediary organizations often sits between platform-facing discourse with everyday work practices. Through training programs, guidelines, and managerial communication, these organizations curate algorithmic explanations and define credible knowledge for workers. These dynamics foreground questions about how interpretive authority of algorithms is organized and how discursively constructed algorithmic knowledge becomes parts of labor management within institutions.
@@
This paper addresses this gap by examining the role of Multi-Channel Networks (MCNs) in China’s live-streaming industry. MCNs are intermediary organizations that operate between individual streamers and platform companies. They help streamers initiate their channel, improve their content, build their personal brand, and make profits, while also shaping what streamers produce and how they work (Zhang et al., 2024; Zhang and Tong, 2024; Liang and Ji, 2024; Wang and Cao, 2024). While MCNs have been widely studied in western contexts such as YouTube as service providers (Siciliano, 2023; Gardner and Lehnert, 2016), less attention has been paid to their epistemic role in algorithm interpretation. We advance this literature by theorizing MCNs not just as economic intermediaries, but as organizational actors that actively interpret and institutionalize algorithmic knowledge. This shift foregrounds how algorithmic interpretation is produced and circulated at scale. China’s live-streaming industry, then, offers a particularly generative setting for this analysis. Live-streaming in China is large scale, highly industrialized, and labor-intensive (Ye et al., 2023; author = John Peikang Sun et al., 2023), with MCNs exerting strong control over labor and visibility. Despite their waning influence on Western markets (Alexander, 2018), MCNs remain influential in China, shaping how streamers understand and respond to algorithmic systems. Given the informational asymmetries in live-streaming work and the limited access many streamers have to reliable explanations of platform algorithms – including many women from rural areas with limited formal education (Xiao et al., 2024a; Tang et al., 2022), MCNs can possess disproportionate epistemic authority in shaping what counts as credible algorithmic knowledge.
@@
Building on such research potential, this study investigates how Chinese MCNs discursively construct algorithms and integrate these constructions into organizational practices in labor management. Drawing on nine months of ethnographic fieldwork and 44 interviews with live-streamers, former live-streamers, and MCN staff, we ask:
  • RQ1: How do MCNs construct organizational interpretations of algorithms?
@@
  • RQ2: How are these interpretations embedded into everyday labor management practices?
  • RQ3: How do live-streamers react to MCNs’ algorithmic interpretations?
In response to RQ1, We found that MCNs construct organizational interpretations of algorithms through a dual framework. Internally, MCN staff treat algorithms as volatile and uncertain systems, relying on probabilistic reasoning to manage performance and organizational risk. Externally, in their interactions with cooperated streamers, MCNs promote simplified and prescriptive narratives that portray platform algorithms as transparent, fair, and responsive to individual effort.
In response to RQ2, we demonstrate how these internal and external algorithmic interpretations are embedded into everyday labor management practices. MCNs translate algorithmic narratives into training programs, performance rankings, aesthetic and behavioral standards, spatial arrangements, and financial dependencies. Through these mechanisms, algorithmic uncertainty is rendered manageable, and responsibility for success or failure is transmitted onto individual streamers through disciplined labor practices.
In response to RQ3, we show that live-streamers’ reactions to MCNs’ algorithmic interpretations unfold along a patterned trajectory. While many initially comply with MCN guidance as a pragmatic response to uncertainty; others develop doubt as promised outcomes fail to materialize; and some ultimately disengage from MCNs or exit the live-streaming industry altogether. These reactions illustrate how algorithmic interpretations are variably accepted, questioned, and resisted over time.
Thus, this study makes three contributions to CSCW and HCI research:
  • First, it foregrounds intermediary organizations as key sites where organizational algorithmic authority is constructed, influencing how algorithmic knowledge is constituted and operationalized in platform economy.
  • Second, it demonstrates how discursively constructed algorithmic interpretations operate as instruments of organizational labor management.
  • Third, it offers design and policy implications by highlighting the role of intermediary organizations in mediating algorithmic authority and governing algorithmic labor.
@@ 2. Background and Related Work
First, we examine related research on algorithm interpretations, such as folk theories, mental models, algorithmic imaginary, and discursive construction of algorithms (subsection 2.1). Second, we analyze the shift from direct algorithmic management (e.g., task allocation or performance metrics) to the softer and more affective forms of control (subsection 2.2). Third, we focus on reviewing literature on Multi-Channel Networks (MCNs) and surrounding studies on the live-streaming industry (subsection 2.3).
2.1. Interpretive Frameworks of Algorithms
This section reviews research that examines how algorithms are understood in everyday practice. Across HCI, CSCW, and digital studies, scholars have developed a range of conceptual frameworks to capture these interpretive processes, including folk theories of algorithms, mental models, algorithmic imaginaries, and recent work on the discursive construction of algorithms. These frameworks foreground how algorithmic systems acquire meaning, shaping how users orient their actions under conditions of algorithmic opacity and uncertainty.
2.1.1. User-Centered Interpretations of Algorithms
A substantial body of research has examined how workers develop practical understandings of algorithmic systems through experience. Within this literature, several closely related concepts have been used to capture different dimensions of user-centered interpretation. Folk theories of algorithms describe the informal, experience-based explanations that users develop to account for algorithmic behavior, often drawing on trial and error, anecdotal evidence, and peer exchange (DeVito et al., 2017; Dogruel, 2021; Huang et al., 2022). Mental models emphasize the internal representations users form about how algorithmic systems function, highlighting how these cognitive structures guide expectations and interaction with complex systems (Ngo et al., 2020). Algorithmic imaginaries foreground the affective, cultural, and symbolic dimensions of algorithmic understanding, attending to how algorithms are imagined within broader social meanings (Bucher, 2018).
Empirical studies show that such understandings often circulate within peer networks, where workers exchange interpretations, compare outcomes, and refine shared explanations of algorithmic behavior (McDaid et al., 2023; Yao et al., 2021; Christin, 2020; Poon et al., 2021). Through these processes, interpretations of algorithms become stabilized and inform everyday decisions about what to produce, when to work, and how to respond to algorithmic feedback (Bonini and Treré, 2024; Lei, 2021). Research on platform labor further demonstrates how these interpretive resources guide routine work practices across domains such as content creation, moderation, and ride-hailing, even when the underlying explanations remain speculative (Bishop, 2019, 2020; Cameron et al., 2021b; DeVito, 2022).
Recent studies also highlight the utility of these interpretive frameworks. For example, Mayworm et al. (Mayworm et al., 2024) show how users mobilize shared explanations of content moderation algorithms as diagnostic tools, using them to adjust strategies, anticipate enforcement, and decide when disengagement becomes necessary. Across these accounts, user-centered interpretations of algorithms function as practical resources that render opaque systems actionable in everyday work.
2.1.2. Discursive Construction of Algorithms and Institutional Mediation
Alongside user-centered and labor-focused approaches, an expanding body of scholarship examines how algorithmic understandings are shaped through institutional discourse and communicative practices. Situated within the broader tradition of algorithmic imaginaries, these studies attend to the ways algorithms are publicly articulated and explained through policy statements, business transparency reports, and educational materials produced by platforms, designers, and regulatory institutions (Ananny and Crawford, 2018; Cellard, 2022; Schulz, 2023). Rather than focusing on individual sensemaking alone, this line of work traces how authoritative accounts of algorithms are assembled and circulated through formal channels, contributing to shared expectations about how algorithms function and what forms of behavior they reward or penalize.
A key theoretical contribution to this literature is offered by Lin ((2025)), who introduces the concept of the “discursive construction of algorithms” to systematize these institutional processes. Lin (Lin, 2025) conceptualizes algorithms as objects whose meanings are actively produced through discourse. As institutional actors increasingly publish accounts of algorithmic operations, algorithms become legible through standardized vocabularies that foreground certain principles, causal logics, and evaluative criteria. In this formulation, algorithms are understood as “dynamic, contextual constructs shaped through discourse (Lin, 2025),” reflecting the situated and contingent ways their meanings are stabilized through communication.
This perspective directs attention to how discursive practices establish interpretive boundaries around algorithms. Lin (Lin, 2025) summarized that, through selective emphasis, simplification, and abstraction, institutional accounts intentionally highlight particular dimensions of algorithmic systems while leaving others opaque or indeterminate. These processes contribute to the normalization of specific interpretations, shaping what is treated as common sense knowledge about algorithms and which forms of uncertainty are rendered acceptable for users.
Emerging research extends this analysis beyond platform companies to examine how discursive constructions of algorithms travel through intermediary actors that connect institutional narratives with everyday labor practices. Studies of MCNs, fan communities, and other third-party organizations illustrate how algorithmic explanations are translated into optimization advice and even collective action frameworks that orient user behavior (Siciliano, 2023; Xiao et al., 2024b). At the same time, existing studies provide limited empirical insight into how intermediary organizations integrate discursively constructed algorithmic accounts into labor management.